OpenStudy (anonymous):

Select the appropriate restrictions for the variables. 6a+12 / a+2 A. a≠–2 B. a≠2 C. a≠6 D. There are no restrictions on the values of variable a.

OpenStudy (theeric):

So, I'm guessing the problem is like this: We want \(\dfrac{6a+12 }{ a+2}\) to be an existing number. But denominators can't be \(0\), because just try and divide something by \(0\). It makes no sense. So it's something you can write down, but it's meaningless mathematically. So we need to make sure that the denominator isn't \(0\). That variable \(a\) can ruin it for us. It can make the denominator be \(0\). Buuuuuuut, if we "restrict" it, so it can be some numbers and not others, then we can say that your fraction there is okay. Like, if \(a\) is going to make the denominator zero, we simply say, "hey, \(a\), you can't be that value that makes the denominator \(0\)!".

OpenStudy (theeric):

So do you have any thoughts on this problem?

OpenStudy (anonymous):

Um...not really, I'm still kinda confused :/ Am I supposed to find the number that's a? or what a represents?

OpenStudy (theeric):

Well, what you need to do, getting straight to the point, is find what \(a\) can be to make the denominator \(0\). Then you want to restrict \(a\) by saying it cannot be that value that makes the denominator zero. So we look at \(\dfrac{6a+12 }{ a+2}\) The denominator is then \(a+2\), right?

OpenStudy (anonymous):

ya

OpenStudy (theeric):

Okay, so what \(a\) will make it \(0\)? Here's a good way to do it. Suppose that the denominator \(\sf does\) equal \(0\), and then solve for \(a\). So, \(a+2=0\implies a=?\)

OpenStudy (anonymous):

2? Because can't you replace the 0 with 2? :/

OpenStudy (theeric):

Are you solving for \(a\)?

OpenStudy (anonymous):

I dont really know....I'm kinda confused on how to solve for a, am I supposed to use the top numbers?

OpenStudy (theeric):

Nope! Just \(a+2=0\). Are you taking algebra?

OpenStudy (anonymous):

Yeah

OpenStudy (theeric):

Okay! So, you know that you want to get \(a\) to its own side of the equation. But you see that you're adding \(2\) to it. How do you get rid of a \(2\) that's being added?

OpenStudy (anonymous):

um..You can use the 6 and 12 I think..sorry I'm really confused :/

OpenStudy (theeric):

Well, forget about the entire problem. .... Now, think about this problem: \(a+2=0\) Find \(a\).

OpenStudy (theeric):

Can you find the answer to that problem?

OpenStudy (anonymous):

Um well, If you can use negatives then I kinda think it would be -2..

OpenStudy (theeric):

Looks good! \(\ \ \ a+2=0\) \(\ \downarrow\ \ \ \ \downarrow\ \ \ \downarrow\) \(-2+2=0\)

OpenStudy (anonymous):

so that's right..?

OpenStudy (theeric):

So, think about this. When \(\color{blue}{a=-2}\), then \(\color{red}{a+2=0}\). So, when you have \(\dfrac{6a+12 }{\color{red}{ a+2}}\) and \(\color{blue}{a=-2}\) then \(\dfrac{6a+12 }{\color{red}{ \color{blue}{-2}+2}}\) \(=\dfrac{6a+12 }{0}\) which is really not a number. So we can't have \(\color{blue}{a=-2}\). So we \(\sf restrict\) \(a\) to not be \(2\). So the answer is \(a\neq2\) (is our restriction)

OpenStudy (anonymous):

ohh ok, so I could have just said 2? instead of -2?

OpenStudy (theeric):

You have to say that \(a\) CANNOT be \(2\). So you say it like \(a\) does not equal \(2\) by writing \(a\neq2\)

OpenStudy (theeric):

I mean negative 2!

OpenStudy (theeric):

So, correction: You have to say that \(a\) CANNOT be \(-2\). So you say it like \(a\) does not equal \(-2\) by writing \(a\neq-2\)

OpenStudy (anonymous):

ohh ok Thank you! :D

OpenStudy (theeric):

If \(a\) does equal \(2\), then things are bad because then \(\quad\dfrac{6a+12 }{ a+2}\) is not a number anymore. It's "undefined." It makes no sense! You're welcome!

OpenStudy (anonymous):

Thank you :)

OpenStudy (theeric):

You're welcome! :)